(a) Find the slope of the tangent line to the graph of the polar curve r = 1 + 2 cos θ at the point where θ = π/3 .

(b) What are the x, y coordinates of the point in the curve r = 1 + 2 cos θ where θ = π/4.
Answer #1

r = 1+2cos\theta

x = cos\theta+2cos^2\theta \; \: y = sin\theta+2cos\theta sin\theta

x = \frac{1}{2}+2\frac{1}{2}\frac{1}{2} \; \: y = \frac{\sqrt{3}}{2}+2\frac{1}{2} \frac{\sqrt{3}}{2}

x =1 \; \: y = \sqrt{3}

x' = -sin\theta-4cos\theta sin\theta \; \: y' = cos\theta+2cos2\theta

x' = -\frac{\sqrt{3}}{2}-4\frac{\sqrt{3}}{2}\frac{1}{2} \; \: y' = \frac{1}{2}-2\frac{1}{2}

x' = -\frac{3\sqrt{3}}{2} \; \: y' = \frac{-1}{2}

dy/dx = \frac{1}{3\sqrt{3}}

Tangent is

y-\sqrt{3} = \frac{1}{3\sqrt{3}}(x-1)

3\sqrt{3}y-9 = x-1

3\sqrt{3}y= x+8

r = 1+ 2\frac{1}{\sqrt{2}} = 1+\sqrt{2}

cos\theta = sin\theta = \frac{1}{\sqrt{2}}

x = y = \frac{1}{\sqrt{2}} (1+\sqrt{2}) = \frac{1}{\sqrt{2}} +1